OpenStudy (anonymous):

the express 2logx - (3logy + logz) is equivalent to...

OpenStudy (watchmath):

\(log x^2-\log y^3-\log z\) \[\displaystyle \log(\frac{x^2}{y^3z})\]

OpenStudy (anonymous):

log\[\log(x^{2}/y ^{3}z)\]